(2x+3y) is divisible by 17,then for what value of k; (9x+ky) is divisible by 17.
Please give the solution.

Grade:11

3 Answers

Shrey
10 years ago
Please reply fast
mycroft holmes
10 years ago
4(2x+3y)+(9x+5y) = 17(x+y). Since RHS is divisible by 17, and so is 4(2x+3y), we must have 9x+5y also divisible. Hence for any k such that k = 5+19m for any integer m, we have 9x+ky divisible by 17
Sher Mohammad IIT Delhi
9 years ago
gcd(2x+ 3y, 9x + ky ) = 17
or gcd(2x + 3y, (9x + ky) - 4(2x + 3y)) = 17
or gcd(2x + 3y, (x + (k-12) y) = 17
gcd ( 2x+ 3y - 2(x + k- 12y), (x + (k-12) y)) = 17
so 27y -2k y is divisible by 17
so 27 - 2k is divisible by 17 or k = 5